MANGO INC.
CONSOLIDATED BALANCE SHEET
September 30, 2017
(dollars in millions)

ASSETS

Current assets:

Cash

$

14,054

Short-term investments

11,401

Accounts receivable

17,718

Inventories

2,138

Other current assets

24,186

Total current assets

69,497

Long-term investments

131,998

Property, plant, and equipment, net

20,915

Other noncurrent assets

12,702

Total assets

$

235,112

LIABILITIES AND STOCKHOLDERS’ EQUITY

Current Liabilities:

Accounts payable

$

30,625

Accrued expenses

18,717

Unearned revenue

8,617

Short-term notes payable

6,398

Total current liabilities

64,357

Long-term debt

29,404

Other noncurrent liabilities

28,253

Total liabilities

122,014

Stockholders’ equity:

Common stock ($0.00001 per value)

1

Additional paid-in capital

25,512

Retained earnings

87,585

Total stockholders’ equity

113,098

Total liabilities and shareholders' equity

$

235,112

Assume that the following transactions (in millions) occurred during the next fiscal year (ending on September 29, 2018):

  1. Borrowed $18,304 from banks due in two years.
  2. Purchased additional investments for $24,500 cash; one-fifth were long term and the rest were short term.
  3. Purchased property, plant, and equipment; paid $9,610 in cash and signed a short-term note for $1,448.
  4. Issued additional shares of common stock for $1,507 in cash; total par value was $1 and the rest was in excess of par value.
  5. Sold short-term investments costing $19,045 for $19,045 cash.
  6. Declared $11,163 in dividends to be paid at the beginning of the next fiscal year.

2. Post each transaction to the appropriate T-accounts (Enter your answers in millions.)Cash 14,054 Short-term Investments Beg. Bal. 11,401 Beg. Bal. End. Bal. 11,401 End. Bal. 14,054 Accounts Receivable Beg. Bal.
Answer #1

Credit Date Date Date Credit Mango Inc. T Accounts Date Debit Date Cash Op bal $14,054 b a $ 18,304 C d $1,507 e $ 19,045 Deb

Mango Inc.
Journal entries
Date Account Debit Credit Calculation
a Cash $     18,304
Long-term debt $     18,304
b Long-term investments $        4,900 =24,500*1/5
Short-term investments $     19,600 =24,500-4,900
Cash $     24,500
c Property, plant, and equipment, net $     11,058
Cash $        9,610
Short-term notes payable $        1,448
d Cash $        1,507
Common Stock $                1
Additional paid-in capital $        1,506 =1,507-1
e Cash $     19,045
Short-term investments $     19,045
f Retained Earnings $     11,163
Dividends Payable $     11,163